There are 4 eight cards, and 12 face cards including the Jack, Queen and King.
So 4 +12 =16, which is the number of eight and face cards you have in a standard deck.
To find the probability, we will have 16/52 = 0.308
Yes if people sin or lie they ecsecute them by chopping their head. Then somone in perticular takes the head and they cut through the head and take out the flesh . Then they sew the cut up and take the eyes out and sew the eyes shut. After to shrink the head they put it in herbs and some other stuff until shrunken . Finally theyput decorations on their head and hair but it is illegal in the United States .
I DO NOT SHRINK HEADS!
The risk of an accident - fatal or otherwise - is not constant over a person's lifetime. Young and inexperienced drivers or elderly drivers with slower reflexes, for example, are at higher risk. Also, different age groups make different numbers of road trips. In real life, therefore, it is quite unreasonable to extrapolate the 1 fatal accident in 4 million trips to 50,000 trips.
If you roll a number cube 100 times about how many times can you expect to roll a 2 explain?
If this is a homework assignment, please consider trying to answer it yourself first, otherwise the value of the reinforcement of the lesson offered by the assignment will be lost on you.
If a number cube (die) contains the numbers 1, 2, 3, 4, 5, and 6, and the cube is fair, then the probability of rolling a 2 is 1 in 6. If you roll the cube 100 times, you would expect to get 2's 100 / 6, or 17 times. However, 100 trials is not a lot of trials, so the experimental outcome might not match the theoretical probability.
The answer is False
What is probability of drawing a royal flush from a pinochle deck?
this is for 5-card poker
(4*2^5) which gives you 128, that's your m
devide it by n, which is the possibilities of hand draws (8C5)
so it will be 128/1712304
and that's your answer.00007547
Three coins are flip at the same time What is the probability that the outcome is at least 2 heads?
There are four outcomes possible (not considering order)
In a standard deck of 52 cards how many are 5 or nines?
There are four 5 and four 9 cards. Thus 8 out of 52 cards are 5 or 9.
In all there are 12 ! ways. If Mr and Mrs S are treated as a single unit, then there are 11! ways of arranging the 10 other people with Mr and Mrs S together.
But these can be Mr-Mrs or Mrs-Mr so, in fact, there are 2*11! ways in which Mr and Mrs S are together.
So the number of ways in which they are NOT together is 12! - 2*11!
= 399,168,000
In the standard normal distribution the means is always?
0
This may not make sense, but it takes a concept known as z-scores into account, which basically standardizes everything.
if repeating is allowed than 1/263= 1/17576, if repeating isn't allowed than it's 1/(26*25*24)=1/15600.
Rolling three dice what is the probability of rolling an odd number?
The probability that, when rolling three dice none will come up odd is broken down by the probability that on odd roll will occur as an independent event. Rolling the first die has a probability of 3/6 that it will be odd. Rolling the second has the probability of 3/6 that it will be odd. Rolling the third has a probability of 3/6 will be odd. Rolling three and coming up with odd is really the probability that no odd numbers occurred, so therefore an even number DID happen (3/6), and an even number DID happen (3/6), and an even number DID happen (3/6). That's really just 1 - (3/6^3) or simplified 1 - (1/2^3). So therefore the probability that an odd number occurred in 3 die rolls is about 87.5%.
What does probability mean in math terms?
you are finding the probable answer in a question
EXAMPLE: Michelle has 10 cards. She is giving 2 to Shannon. What would the fraction, the dot (decimal) and the percent be?
ANSWER: The fraction is... 2-10
The decimal is... .20
The percent is...20%
What is the theoretical probability of throwing a coin and it landing on heads?
Theoretical probability is the number of ways something can occur divided by the total number of outcomes. So, the theoretical probability of throwing a coin and it landing on heads is 1/2 or 0.5 or 50%.
A coin is tossed two times what is the probability of obtaining at least one tail?
3/4
The only way a tail doesn't show up is if all heads come up.
P(all heads) = (1/2)*(1/2)=1/4
P(at least one tail)=1-P(all heads)=1-(1/4)=3/4
What events occurred on December 7th?
Japanese bomb Pearl Harbor, Hawaii; also attack the Philippines, Wake Island, Guam, Malaya, Thailand, Shanghai and Midway.
Consider a coin toss. The probability of the coin coming up head is 1/2, the probability of the coin coming up tails is 1/2. No matter how many times you flip the coin the probability of any particular toss coming up heads or tails is always 1/2.
BUT
If you consider two coins tosses as a set then you have four possible outcomes:
first coin heads - second coin heads, first coin heads - second coin tails,
first coin tails - second coin heads, first coin tails - second coin tails
for ease of demonstration lets denote heads as H and tails as T
the above combination would then be HH,HT,TH,TT
the probability of any one of the combinations would be 1/2 for the first coin and 1/2 for the second coin so all combinations have a probability of 1/4. (1/2x1/2=1/4 I hope you realize.)
But, if we don't care which coin is H or T then the two combinations HT and TH are the same and we combined their probabilities into one 1/4+1/4=1/2
HH,(HT,TH),TT
1/4 1/2 1/4
What are the odds of 2 dice totaling a 7 on one throw?
Answer: 1/6
It does not matter what number the first die lands on: you will always have a 1/6 chance that the second die makes a total of 7: if it lands on a 6 while the first was 1; if it lands on a 5 while the first was 2; and so on.
If a die is tossed once probability of it landing on a 6?
Since there are 6 possible outcomes, and you want the probability of obtaining one of the outcomes (in your case 6), the probability of it landing on a 6 is 1/6.
How many ways can 3 six sided dice be thrown to equal 7?
15..... 1,1,5 1,2,4 1,3,3 1,4,2 1,5,1 2,1,4 2,2,3 2,3,2 2,4,1 3,1,3 3,2,2 3,3,1 4,1,2 4,2,1 5,1,1
put one cup into another cup, then put 5 marbles in each.
Basic arithmetic has been developed by nearly every culture, so there isn't one definite person who discovered it.
Random or running numbers to win lottery?
Random lottery numbers, every time! I think it's great to mix it up, but I also like to see my numbers before I buy my ticket. I play the euro millions a bit more than the lottery as the rollover jackpots always look appealing! In theory it doesn't matter how you select the numbers so long as you pick the rquired number of numbers.
What is the probable ion formed from Br?
Br-. Bromine will grab an electron to make itself more stable, which makes it a negative ion.